### v1

Abstract
Considering the close connection between action recognition and human pose estimation, we design a Collaboratively Self-supervised Video Representation (CSVR) learning framework specific to action recognition by jointly considering generative pose prediction and discriminative context matching as pretext tasks. Specifically, our CSVR consists of three branches: a generative pose prediction branch, a discriminative context matching branch, and a video generating branch. Among them, the first one encodes dynamic motion feature by utilizing Conditional-GAN to predict the human poses of future frames, and the second branch extracts static context features by pulling the representations of clips and compressed key frames from the same video together while pushing apart the pairs from different videos. The third branch is designed to recover the current video frames and predict the future ones, for the purpose of collaboratively improving dynamic motion features and static context features. Extensive experiments demonstrate that our method achieves state-of-the-art performance on the UCF101 and HMDB51 datasets.
keywords:
Video representation learning, Self-supervised learning, Action recognition, Human pose prediction
1 Introduction
Human action recognition is a fundamental task in computer vision and has gained increasing attention in recent years. With a large number of labeled videos, the convolutional neural network (CNN) has achieved great success in human action recognition by learning powerful spatiotemporal representation features [32, 33, 81, 80]. However, the cost of collecting and tagging large amounts of videos is enormous. Fortunately, there is a huge amount of unlabeled videos on the Internet. A common practice is to use a large amount of unsupervised video for robust feature representation learning and then transfer to downstream tasks such as human action recognition, which deserves further exploration.
Self-supervised learning has emerged as a promising approach for unsupervised learning. Its goal is to learn powerful feature representation by designing appropriate pretext tasks which automatically generate free annotations for unlabeled data and transfer the representations to various downstream tasks. This approach has achieved remarkable success in areas such as image recognition, object detection, and image segmentation. Besides the self-supervised learning on images [55, 60], there has been growing interest in exploring self-supervised learning for video representation with the increasing availability of large-scale unlabeled video data.
Early video self-supervised learning approaches mainly focus on effectively exploiting contextual relationships within videos [58, 56]. Over the years, self-supervised methods have become more powerful and diverse can generally be categorized into two groups: discriminative classification / regression methods [94, 86] and generative prediction ones [29, 6, 4]. Discriminative methods focus on learning to distinguish between different samples, while generative methods aim to predict missing information. Recent efforts have been made to learn general video representations that can be applied to a range of downstream tasks. This includes video contrastive learning [94, 18, 69, 1], as well as cross-modal learning [115, 105], which learns from different modalities such as video, audio and text. Although these methods achieve promising results for various video based downstream tasks, they do not explicitly consider designing pretext tasks for action understanding, which may lead to sub-optimal solutions for action recognition. One interesting work, DCM [160] designs a two-stream network to separately learn action-related features, i.e., context and motion features, which achieves impressive results for action recognition. However, DCM fails to jointly consider the learning of context and motion features, lacking comprehensive view of these two features.
In this work, we propose the Collaboratively Self-supervised Video Representation (CSVR) framework, which jointly learns dynamic motion and static context features for the downstream action recognition task. As illustrated in Figure 1, CSVR consists of three main representation learning branches, i.e., pose prediction branch, context matching branch, and a collaborative training branch named video generating branch. To enhance the performance of action recognition, we resort to the strong relationship between human pose and actions. Specifically, we employ a generative pose prediction branch based on a conditional generative adversarial network (CGAN) [83], which predicts future poses by taking current pose sequences as input. By excluding the potential motion noise from backgrounds, our method can effectively learn dynamic motion features highly related to action. Video key frames, known as the intra frame (I-frame) in video compression domain, contain valuable context information. To leverage the abundant pattern in I-frame for action recognition, we incorporate a discriminative context matching branch, which works by pulling the features of clips and I-frames from the same videos closer together, while simultaneously pushing apart pairs from different videos. Finally, we design the collaborative video generating branch to jointly optimize the dynamic and static features. The branch contains two video generating networks, one aims to reconstruct current video clips, another is trained for predicting future video clips. The video generation process requires both static context information and dynamic motion information, which provides a powerful collaborative learning objective, assisting the previous two branches in learning excellent static and dynamic features.
Figure 1: An overview of our self-supervised learning progress. Our CSVR contains three branches: pose prediction branch, context matching branch, and video generating branch. The pose prediction branch uses poses from input video clips to predict future poses. Dynamic motion feature $f_{d}$ is originating from the encoder of pose generator. The context matching branch captures video clip features and intra frame (I-frame) features by CNNs, and then casts a contrastive loss on learned features to pull the feature pairs from the same video clips together. Through the contrastive learning method, video clip feature $f_{s}$ can contain rich static context information. After fusing previous features by the integration layer, integrated feature $f_{i}$ is fed into the video generating networks, try to reconstruct current video clips and predict future video clips, respectively. The video generating branch jointly optimizes all the branches, leading to learn more comprehensive video representation.
Extensive experiments on three popular datasets with various backbones show that our CSVR achieves state-of-the-art performance for two downstream action understanding tasks, i.e., action recognition and video retrieval.
We summarize the contributions in the following:
1) We introduce a novel framework focusing on action recognition named Collaboratively Self-supervised Video Representation learning (CSVR), which simultaneously learns dynamic motion and static context features using three branches with discriminative and generative pretext tasks.
2) Considering the strong connections between human pose and actions, we design a generative pose prediction branch that predicts future poses from current pose sequences, enabling efficient and effective extraction of dynamic motion features.
3) Our carefully designed collaborative video generating branch enables joint optimization of dynamic motion and static context features, resulting in more comprehensive representations that benefit downstream action understanding tasks.
2 Related Work
As our method is primarily focused on self-supervised video representation learning for action recognition, in this section, we provide a brief overview of some typical works in the fields of video action recognition and self-supervised video representation learning. Additionally, we review some works on Generative Adversarial Networks (GAN) [84] as GAN-based generative branches play a crucial role in our method.
2.1 Video-Based Action Recognition
Videos contain motion information with a complex temporal structure, which requires spatio-temporal feature learning. To address this, numerous advanced 3D-CNN architectures have been proposed [32, 33, 150, 152]. Recently, with the inspiration of Transformer[143] architectures in natural language processing, researchers have attempted to extend Vision Transformer (ViT) [89] for video modeling. Many works [4, 90, 91] propose different variants for spatiotemporal representation learning, and adapt to video action recognition task. However, learning favorable motion representations from original video clips remains a challenging task. To this end, various movement information have been previously examined to help extracting human action information from raw video, such as optical flow [85], motion vectors [116, 128], and pose data [19, 20]. With the advancement of body pose estimation methods [79], several approaches have utilized pose sequence information for assisting video action recognition. Previous works directly encode temporal correlation within pose sequences for action understanding [36, 38, 40]. The high efficiency of video networks also leads to breakthroughs in self-supervised video representation learning for action recognition [160, 85, 43, 44]. Some of them [43, 44] use pose information as an attention or supervision signal to mine video representations more suitable for action understanding tasks.
2.2 Self-Supervised Video Representation Learning
Self-supervised video representation learning methods aims to leverage unlabeled video data to learn representations that can be easily transferred to downstream tasks. Existing methods for self-supervised video representation learning can be generally categorized into three types: context-based methods, contrastive learning and generative algorithms[10]. Early context-based methods focus on designing pretext tasks that exploit contextual relationships within videos, such as speed prediction [160], pace prediction [56], and jigsaw solving [58]. Contrastive learning methods focus on how to construct positive and negative examples without annotations and pull positive examples closer to each other than negative ones in training process [68, 25, 156, 92, 88]. Interestingly, in contrast to the CNN-based approaches, there are a few very recent studies [92, 88] leveraging contrastive learning for transformers. Recently, there has been a surge of research into Transformer-based generative mask video modeling methods [6, 87, 93]. These works learn representations by exploiting co-occurrence relationships between video patches in reconstructing process. Moreover, [93] tries to reconstruct features from multiple sources to perform self-supervised pretraining in videos.
Although self-supervised video representation learning methods achieve impressive performance on various downstream tasks including action recognition, most of them are not specially designed for action-oriented representation learning. Differently, DCM [160] and P-HLVC [19] aims to learn generalizable features specific to action understanding. They simply extract motion features by contrastive learning and leverage diverse motion information, i.e., pose and motion vectors as self-supervised signals. Inspired by these approaches, we propose a Collaboratively Self-supervised Video Representation (CSVR) learning framework that simultaneously considers pose prediction and context matching as pretext tasks. By jointly optimizing the whole framework by video generating branch, CSVR provides a more comprehensive feature representation for downstream action recognition task.
2.3 Generative Adversarial Networks
Generative adversarial networks (GAN) [84] usually consist of two encoder-decoder architecture networks: a generator producing fake data and a discriminator distinguishing between fake and real data. However, the inherent randomness of GAN output makes it difficult to control the generated data. To address this issue, Conditional GANs (CGAN) [83] are proposed for incorporating input conditions such as class labels to regulate the synthetic outputs. Recently, more efforts are devoted to video generation tasks. TGAN v2[77] is a video generation network that can efficiently generate realistic videos. Some works [163, 164] prove the feasibility of learning video representation from generative learning methods. We believe the process of generating diverse samples utilizing GANs can better capture human motion and learn representations which benefit the action understanding tasks. In CSVR, we employ 1D-CNN-based CGAN in pose prediction branch and R3D-based TGAN v2 in video generating branch. To enhance the stability and efficiency of our training process, we also adopt the gradient penalty from WGAN-GP [72] and the parameter update strategy from OSGANs [50]. These carefully-designed objectives help our method efficiently generate more accurate and diverse samples, thus improve the quality of learned dynamic motion representations.
3 Methodology
3.1 Overall Framework
In this paper, we propose a novel framework for action recognition called Collaboratively Self-Supervised Video Representation learning (CSVR). The framework aims to learn robust and comprehensive feature representation for action recognition, which consists of both dynamic motion and static context features. To achieve this, we design three branches: a generative pose prediction branch for learning dynamic motion information, a discriminative context matching branch for learning static context representation, and a collaborative video generating branch to jointly optimize all branches. Figure 1 provides an overview of our method.
Human pose modeling is an intuitive and effective way to represent human motion. Therefore, how to effectively learn the dynamic information contained in poses has become the core issue for human action analysis. Previous work has made great progress in action recognition tasks using optical flow [155] and motion vectors [160]. However, these approaches may perform poorly in practical scenarios due to noise information, such as irrelevant background movement and low-quality video noise, which can be naturally relieved by extracting poses from raw videos. Although some works [19] try to learn dynamic information from poses by contrastive learning approaches, which may fail to fully leverage temporal information in the pose sequences, leading to unsatisfactory results in action recognition task. In CSVR, we resort to designing a generative pose prediction branch to capture dynamic information. Specifically, we use OpenPose [79] to estimate poses from raw video clips [^1] Although we use the off-the-shelf pose estimator OpenPose, which is trained with pose supervision, we provide no additional supervision for the action recognition tasks. It does not violate the motivation of the self-supervised learning., and then deploy CGAN to predict future poses from the current ones. To make the training process even more concise, we utilize time-efficient unidimensional convolutional neural networks (1D-CNNs), which output pose sequences all at once.
Besides the dynamic information from human poses, the context information is also important for action recognition. For the video compression, context information is roughly separated into intra frames (I-frame), which contain the main content of the video clip. Thus, we conduct a context matching branch that compares global features from video clips with static image features from I-frames to learn static context information. Specifically, we randomly select video clips and extract I-frames following the practice of [5]. We then utilize the MultiInstance-InfoNCE (MI-InfoNCE) loss [161] to learn static context features by matching the video and I-frames. This branch pulls together features of videos and I-frames from the same video clips while pushing apart pairs from different clips. By matching still images to video clips, the learned representation is supposed to capture the contextual information beneficial for downstream action recognition task.
Since the pose prediction and context matching branches just learn dynamic and static features separately, it is significant to effectively integrate them to obtain comprehensive representations for downstream tasks. The video generating branch includes an integration layer and two video generation networks that provide powerful target for collaborative learning. The integration layer combines the dynamic and static features by Adaptive Instance Normalization (AdaIN) [120]. We then feed the integrated features into TGAN-based [77] video generation networks for two parallel tasks, video reconstruction and video prediction.
Our collaborative training framework leverages the strong correlation between human action and poses, as well as between context and I-frames. By jointly learning dynamic motion and static context features with three branches, rich representations can be achieved that contributes to action-related downstream tasks, such as action recognition. We will introduce the detail of each branch in the following sections.
3.2 Generative Pose Prediction Branch
The generative pose prediction branch is conducted to learn dynamic motion features, which consists of two modules, i.e., a pose generator ($G_{p}$) and a pose discriminator ($D_{p}$), similar to the structure of CGAN [83]. Using the poses from current video clips as input, the objective of pose generator ($G_{p}$) is to generate plausible future poses. The feature from the last convolutional layers of encoder in $G_{p}$ serve as the source of the dynamic motion feature $f_{d}$. The structure of the pose prediction network is illustrated in the upper half of Figure 1.
Suppose $p_{t}\in\mathbb{R}^{2N}$ denotes the human pose captured in the $t$-th video frame, where $N$ is the number of skeleton joints. Given a randomly selected $T$-frame video clip, we apply OpenPose human skeleton detector [79] to estimate the current pose sequence $P_{in}=(p_{0},p_{1},\cdots,p_{T-1})\in\mathbb{R}^{2N}$ and the future pose sequence $P_{gt}=(p_{T},p_{T+1},...,p_{T+T^{\prime}-1})$ of length $T^{\prime}$ frames. The input to $G_{p}$ is a latent noise vector $z_{p}$ and the current pose sequence $P_{in}$. After applying a one-dimensional filter along the time direction, $G_{p}$ produces a sequence of poses for the next $T^{\prime}$ frames, denoted as $P_{gen}=(p^{\prime}_{T},p^{\prime}_{T+1},...,p^{\prime}_{T+T^{\prime}-1})$. During training, $G_{p}$ learns to generate pose sequences that match the target distribution, while $D_{p}$ aims to distinguish real pose sequences $P_{gt}$ from generated ones $P_{gen}$. The objective function is formulated as:
$\begin{split}\min_{G_{p}}&\max_{D_{p}}\mathcal{L}_{pose}=\\ &E_{P_{gt}}[\log D_{p}(P_{gt})]+E_{z}[\log(1-D_{p}(G_{p}(z_{p}|P_{in}))]+\\ &\frac{1}{T^{\prime}}\sum^{T+T^{\prime}-1}_{t=T}||p_{t}-p^{\prime}_{t}||_{2}^{2}+E_{p}[(\parallel\bigtriangledown D_{p}(P|P_{in})\parallel_{2}-1)^{2}]\end{split}$ (1)
The objective function mainly consists of two parts: L2 reconstruction loss contributes to the generation of accurate future poses, and the adversarial loss helps $G_{p}$ to learn the underlying distribution of the input data, leading to more realistic poses. In addition, we introduce the WGAN-GP gradient penalty [72] (denoted as $E_{p}[(\parallel\bigtriangledown D_{p}(P|P_{in})\parallel_{2}-1)^{2}]$ in Equation 1) and OSGANs parameter update strategy [50] to improve training stability and efficiency. Subsequently, we add convolutional layers and MLPs after the last layer in encoder of $G_{p}$ to keep the dimension of dynamic motion feature $f_{d}$ align with static context feature $f_{s}$.
3.3 Discriminative Context Matching Branch
Besides the dynamic motion information, the context information is also significant for human action analysis. Thus, we conduct a discriminative context matching branch to learn static context features in self-supervised manner.
As depicted in the blue box of Figure 1, the V-Network (Video Network) and I-Network (I-frame network) are used to extract video and context features, respectively. We extract features $x_{i}\in\mathbb{R}^{C_{1}\times T_{1}\times H_{1}\times W_{1}}$ from a random video clip $i$ and feature $z_{i}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ from the corresponding I-frame of the same video, where $C_{1}$, $T_{1}$, $H_{1}$, and $W_{1}$ denote the channel, time, height, and width of video clips, respectively. We use global average pooling to obtain their global representations $x_{i}^{pool}\in\mathbb{R}^{C_{1}}$ and $z_{i}^{pool}\in\mathbb{R}^{C_{2}}$ We then apply two-layer MLP heads, $g^{V}$ and $g^{I}$, by taking the features of the video clip and I-frame as inputs, i.e., $x_{i}^{*}=g^{V}(x_{i}^{pool})\in\mathbb{R}^{C}$ and $z_{i}^{*}=g^{I}(z_{i}^{pool})\in\mathbb{R}^{C}$. Finally, the Multi-Instance-InfoNCE (MI-InfoNCE) loss [161] is applied to pull the video clip and I-frame feature pairs from the same video closer and push pairs from different video further away:
$\begin{split}\mathcal{L}_{context}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\sum\exp(\frac{\cos(z^{*}_{i},x^{*}_{i})}{\tau})}{\sum\exp(\frac{\cos(z^{*}_{i},x^{*}_{i})}{\tau})+\sum_{k\neq i}^{B}\exp(\frac{\cos(z^{*}_{k},x^{*}_{i})}{\tau})}\end{split}$ (2)
where $z_{i}^{*}$ and $z_{k}^{*}$ denote the positive and negative I-frame features, $B$ denotes the number of samples in the minibatch, $\cos(z,x)=(z^{T}x)/(||z||_{2}\cdot||x||_{2})$ denotes the cosine similarity between $z$ and $x$, and $\tau$ is a temperature adjusting the scale of cosine similarities. Considering that there may be multiple I-frames in one video clip, we set all I-frames in video clip $i$ as positive samples. Alternatively, the negative samples are I-frames from other videos in the same minibatch, denoted as $z^{*}_{k}$ where $k\neq i$. Different from original InfoNCE loss [118], MI-InfoNCE loss enlarges the quantity of positive samples in each minibatch, which is more appropriate for the current situation. After pretraining on large video dataset, the video feature $x_{i}^{*}$ from the last layer of V-Network will serve as the static context feature $f_{s}\in\mathbb{R}^{C}$ in our model.
3.4 Collaborative Video Generating Branch
After achieving the dynamic motion feature $f_{d}$ and tatic context feature $f_{s}$, we further conduct a collaborative video generating branch based on TGAN v2 [77] to learn more comprehensive video features by fusing these two features. This branch consists of three main components: a feature integration layer, a video reconstruction network and a video prediction network.
Feature Integration Layer. Inspired by the prior work [120, 119], we adopt Adaptive Instance Normalization (AdaIN) for feature fusion. We firstly align the dimensions of the dynamic motion feature $f_{d}$ and static context feature $f_{s}$ by MLPs. The fusion of $f_{d}$ and $f_{s}$ is accomplished through a denormalization process. The detailed structure of the feature integration layer is shown in Fig. 2.
Specifically, we perform batch normalization [51] on the static context feature $f_{s}$:
$\begin{split}\bar{f_{s}}=\frac{f_{s}-\mu}{\sigma}\end{split}$ (3)
where $\mu$ and $\sigma$ are the means and standard deviations of the channel-wise activations within an input minibatch. Then, we design two parallel branches from $\bar{f_{s}}$ for dynamic motion integration and adaptive attention mask.
For dynamic motion integration, we compute the factor $D$ by denormalizing the normalized $\bar{f_{s}}$, formulated as:
$\begin{split}D=\gamma_{d}\otimes\bar{f_{s}}+\beta_{d}\end{split}$ (4)
where $\gamma_{d}$ and $\beta_{d}$ are modulation parameters obtained through convolution of the dynamic motion feature $f_{d}$. They share the same tensor dimensions with $\bar{f_{s}}$.
The core issue of the integration layer is how to adaptively adjust the integration ratio of dynamic motion feature and static context feature to collaboratively facilitate network training. Therefore, we introduce an attention mechanism to the integration layer that dynamically adjusts the weights of $f_{d}$ and $f_{s}$ during the video reconstruction process. Specifically, we generate an attention mask $M$ from $\bar{f_{s}}$ and normalize the mask via a Sigmoid function. Finally, the integrated feature $f_{i}$ is obtained by an element-wise combination of $D$ and $f_{s}$, weighted by the mask $M$, which is formulated as:
$\begin{split}f_{i}=(1-M)\otimes D+M\otimes f_{s}\end{split}$ (5)
Afterwards, the integrated feature $f_{i}$ is taken as the input of video generation networks. The following figure illustrates the detailed structure of integration layer.
Figure 2: Detailed structure of integration layer. In the process of generating the integrated feature $f_{i}$, we use the parameters generated from $f_{d}$ to denormalize the normalized $\bar{f_{s}}$, and combine the denormalized feature $D$ and static context feature $f_{s}$ with the an attention mask $M$.
Video Generation Networks. We employ the video generating branch to learn dynamic motion feature and static context feature collaboratively. Our framework adopts TGAN v2, a time-efficient video generation network based on CGAN.
In the green box of Figure 1, the video generating branch contains two similar TGANs, named video predictor and video reconstructor. The video predictor consists of a generator and a discriminator, denoted as $G^{p}_{v}$ and $D^{p}_{v}$, respectively. Taken the integrated feature $f_{i}$ and latent noise $z^{p}_{v}$ as input, the video predictor tries to generate future video clips. The video reconstructor has a similar structure, taking $f_{i}$ and latent noise $z^{r}_{v}$ as input, seeks to reconstruct current video clips simultaneously. As shown in the upper half of the green box in 1, the generator and discriminator of video reconstructor are denoted as $G^{r}_{v}$ and $D^{r}_{v}$, respectively. We utilize adversarial loss and L2 loss for each TGAN, where the L2 reconstruction loss calculates the difference between the generated sample and the ground truth at the pixel level. The loss objective for video generating branch $\mathcal{L}_{video}$ involve two objectives $\mathcal{L}_{r}$ and $\mathcal{L}_{p}$:
$\begin{split}\min_{G^{r}_{v}}&\max_{D^{r}_{v}}\mathcal{L}_{r}=\\ &E_{V_{r}}[\log D^{r}_{v}(V_{r})]+E_{z}[\log(1-D^{r}_{V}(G^{r}_{V}(z^{r}_{v}|f_{i}))]+\\ &\frac{1}{THW}\sum^{THW}||v_{t}^{r}(h,w)-g_{t}^{r}(h,w)||_{2}^{2}\end{split}$ (6)
$\begin{split}\min_{G^{p}_{v}}&\max_{D^{p}_{v}}\mathcal{L}_{p}=\\ &E_{V_{p}}[\log D^{p}_{v}(V_{p})]+E_{z}[\log(1-D^{p}_{V}(G^{p}_{V}(z_{v}^{p}|f_{i}))]+\\ &\frac{1}{THW}\sum^{THW}||v_{t}^{p}(h,w)-g_{t}^{p}(h,w)||_{2}^{2}\end{split}$ (7)
$\begin{split}\mathcal{L}_{video}=(\mathcal{L}_{r}+\mathcal{L}_{p})/2\end{split}$ (8)
where $V_{r}$ and $V_{p}$ denote the generation target of each generator videos and $T$, $H$, $W$ represent the temporal and spatial dimensions of video clip, respectively. $v_{t}^{r}(h,w)$ and $v_{t}^{p}(h,w)$ indicates the pixel value at the $(h,w)$ coordinates on the $t$-th frame in the ground truth video. The generated videos are denoted as $g_{t}^{r}(h,w)$ and $g_{t}^{p}(h,w)$. Similar to pose prediction branch, we use the OSGANs parameter updating strategy [50] to simplify the training process. Two parallel training objectives enhance the understanding of the temporal relationship of the video, leading to a more comprehensive video representation.
3.5 Joint Optimization
In our CSVR, we jointly optimize the three branches by linearly combining the losses from each branch:
$\begin{split}\mathcal{L}=\sigma_{p}\mathcal{L}_{pose}+\sigma_{c}\mathcal{L}_{context}+\mathcal{L}_{video}\end{split}$ (9)
where the $\sigma_{p},\sigma_{c}$ are scalar hyper-parameters. All the networks in three branches are jointly optimized with $\mathcal{L}$ in an end-to-end fashion. Finally, for fine-tuning on downstream tasks, we utilize the pretrained parameters from the V-Network as the initialization.
4 Experiments
In this section, we conduct experiments on three popular action recognition datasets, i.e., UCF101 [121], HMDB51 [122], and Kinetics400[148] to verify the effectiveness of our CSVR. Besides, we also conduct ablation studies to investigate the significance of each branch in our method.
4.1 Dataset
The experiments of CSVR are mainly conducted on three datasets. UCF101 contains 13,320 videos covering 101 action classes, while HMDB51 consists of 6,766 videos in 51 classes. Following previous methods [94, 160], we use split 1 of UCF101 and HMDB51 in our experiments. Kinetics400 is a large-scale dataset containing 246K videos in its training set which includes 400 classes of human actions. During the experiments, we pretrain our CSVR on UCF101 training set and Kinetics400 training set, respectively. Then we finetune the pretrained V-Network to two downstream tasks, i.e., action recognition and video retrieval, on UCF101 and HMDB51, respectively.
4.2 Implementation Details
Backbone For context matching branch, we explore four 3D-CNN backbones as V-Network, i.e., C3D [149], R(2+1)D[150], R3D [151, 80], and S3D [152] to extract video representations. Moreover, we adopt R2D-10 as I-Network to extract features from I-frame. For pose prediction branch, we utilize 1D-CNN-based CGAN for pose prediction. In addition, R3D-based TGAN v2 is leveraged in video generating branch. It should be noted that during inference, the major computation cost is from video representation extraction V-Network which is mainly determined by various 3D-CNN backbones.
Pretraining Since our CSVR consists of three branches, we firstly pretrain each branch separately and then jointly optimize them in self-supervised manner for stable training. Firstly, CGAN in pose prediction branch and TGAN v2 in video generating branch are pretrained with respective training objectives in unsupervised manner. Then the whole CSVR network is pretrained in self-supervised manner, aiming for a video network with comprehensive feature. During self-supervised learning progress, integrated feature $f_{i}$ is utilized as a conditional variable to drive the generators in producing diverse videos. It is worth noting that GANs are not frozen during self-supervised pretraining.
For self-supervised pretraining, we randomly crop each input video clip to 16 frames with a temporal stride of 4 and resize them to 112$\times$112. We use a batch size $B$ of 16 and set the learning rate to $0.0005\times B$. We use SGD as the optimizer, setting the weight decay to 0.005 and the momentum to 0.9. In each training step of context matching branch, we sample all the I-frames as positive samples and three negative frames from other video. Data augmentation approaches, i.e., random crop, random flip, Gaussian blur, and color jitter are conducted on video clips and I-frames for better performance.
As for the generation of $f_{i}$, the feature of the last layer in pose generator is utilized as the the dynamic motion feature $f_{d}$. After investigating the selection of hyper parameters, which is introduced in Section 4.3, we set $\sigma_{p}=\sigma_{c}=0.75$ in our final experiment.
Finetuning After pretraining our CSVR on video dataset, the integration feature $f_{i}$ from the V-Network is utilized to finetune on downstream tasks. We conduct experiments on two widely-used human action datasets, i.e., UCF101 and HMDB51, and report Top-1 accuracy of action recognition task and Recall at k (R@k) of video retrieval task on each dataset.
For action recognition task, we initialize the backbone with the pretrained parameters except for the last fully connected layer. We finetune the model with a batch size $B$ of 32 and use the cosine annealing scheduler to decay the learning rate during training. For UCF101, we set the learning rate to $lr=0.0001\times B$ and the weight decay of the SGD optimizer to 0.003. For HMDB51, we set the learning rate to $lr=0.0002\times B$ and the weight decay to 0.002. We incorporate a dropout ratio of 0.3 in the second-to-last layer of the network during finetuning. Data augmentation methods, e.g., random crop, random flip, Gaussian blur, and color jitter are applied to increase the diversity of the training data.
For the video retrieval task, we sample 10 video clips with a sliding window and use the average of their global features as the representation of the video. The representations of videos from the test set are utilized to query the k-nearest neighbours in the training set, and Recall at k (R@k) is utilized for the evaluation metric. The network settings, e.g., batch size, learning rate and dropout ratio, are kept the same as the action recognition task. During retrieval, we directly use the representation from the pretrained video network for evaluation. For this task, we do not apply any augmentation except for random crop.
4.3 Ablation Study
To further investigate the effectiveness of each branch of CSVR, we conduct experiments by firstly pretraining on UCF101 training set and then finetuning on the same set for action recognition. We choose R(2+1)D as the backbone of V-network for all ablation study experiments.
Table 1: Action recognition performance comparison of different branches. 4-th row represents the integrated feature $f_{i}$ without optimization from video generating branch, while in the last row, $f_{i}$ is learned by the full CSVR pretraining framework.
$f_{d}$ $f_{s}$ $f_{i}$ $V$ UCF101 HMDB51
(Training from scratch) 65.0 32.5
75.1 38.4
73.1 40.2
81.0 46.3
90.3 56.5
Table 2: Correlation between pretext and downstream tasks. The linear relationship between the results indicates the effectiveness of our pretraining. MSE score is increased by a factor of 100 to emphasize the change. All the experiments are conducted on UCF101 dataset.
Task (Evaluation) Pretraining epoch
1 5 10 20 50 120
[…]
Effectiveness of Different Branches. Since our method includes three different branches, we evaluate the contribution of each branch to the performance improvement. Table 1 shows the final action recognition performance on UCF101 and HMDB51 with features from various branches as the initialization for finetuning on downstream action recognition. For inference, we transfer different latent features to V-Network and finetune them on downstream tasks. As seen, both the static context feature $f_{s}$ and dynamic motion feature $f_{d}$ significantly improve the action recognition accuracy compared to the baseline of training from scratch, which demonstrate the superiority of designing context matching and pose prediction branches. Interestingly, the dynamic motion feature $f_{d}$ achieves better results than static context feature $f_{s}$ on HMDB51, while $f_{d}$ underperform $f_{s}$ on UCF101. This may be because most classes in UCF101 can be classified by context information, while classes in HMDB51 are mainly distinguished by the motion information. Besides, we achieve the integrated feature $f_{i}$ by fusing the static context feature $f_{s}$ and dynamic motion feature $f_{d}$ with AdaIN. By employing $f_{i}$ for the downstream task finetuning, the improvement on accuracy further illustrates the complementarity of static context and dynamic motion features for action recognition. Finally, we introduce video generating branch to further improve the integrated feature $f_{i}$ by jointly optimizing three branches. The accuracy on UCF101 and HMDB51 is significantly improved to 90.3% and 56.5%, respectively, which shows the effectiveness of collaboratively learning static context and dynamic motion features for action recognition.
Correlation between pretext and downstream tasks. We conduct experiments to explore the correlation between the pretext tasks and the downstream task by measuring their performance during different pretraining epochs. The performance measurements include the top-1 accuracy of context matching, mean squared error (MSE) of pose prediction, inception score (IS) of generated videos and top-1 accuracy of downstream action recognition. The results in Table 2 show a consistent improvement in action recognition accuracy as the performance of the pretext tasks increases, which indicates the strong correlations between various pretext and downstream tasks.
Table 3: Comparison of $f_{d}$ from different layers of $G_{p}$ encoder.
Table 4: Comparison of $f_{s}$ with different loss functions.
Layer UCF101 HMDB51
Layer 2 72.0 36.9
Layer 4 74.5 37.2
Layer 6 75.1 38.4
Objective UCF101 HMDB51
L2 Loss 70.2 33.4
InfoNCE 72.5 40.2
MI-InfoNCE 73.1 40.2
Table 4: Comparison of $f_{s}$ with different loss functions.
Selection of Dynamic Motion Feature. In order to locate the hidden layer which enriches the most dynamic information, we pretrain the pose generator $G_{p}$ on poses from UCF101 dataset and use features from different layers of encoder in $G_{p}$ as dynamic motion features $f_{d}$. We report the results of different $f_{d}$s finetuned on action recognition task. As shown in Table 4, $f_{d}$ from the last layer achieves the best performance.
Selection of Contrastive Loss. As shown in the lower half of Figure 1, we employ a contrastive loss between video feature and I-frame feature. The widely used InfoNCE loss [118] compares the feature with limited positive samples and multiple negative samples, which can be formalized as:
$\begin{split}\mathcal{L}_{context}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\frac{\cos(z^{*}_{i},x_{i})}{\tau})}{\exp(\frac{\cos(z^{*}_{i},x_{i})}{\tau})+\sum_{k\neq i}^{B}\exp(\frac{\cos(z^{*}_{k},x_{i})}{\tau})}\end{split}$ (10)
where $x_{i}$ denotes the global video representation, and $z_{i}^{*}$ and $z_{k}^{*}$ denote the positive and negative I-frame features, other symbols are the same as Equation 2. However, there may exist multiple I-frames in a single video clip, indicating more positive samples. Hence, we apply more advanced Multi-Instance-InfoNCE (MI-InfoNCE) [161] as the contrast loss in static context branch. Comparing to the traditional InfoNCE loss utilized in [160], MI-InfoNCE loss enables more adequate comparison between positive and negative samples, resulting in better performance on downstream tasks. To further verify the effectiveness of contrastive loss, we also conduct ablation experiments with L2 loss. When calculating InfoNCE and L2 loss, we randomly select an I-frame from the video clip as learning objective to ensure that the loss function produces correct outputs. As shown in Table 4, MI-InfoNCE achieve optimal results among three losses. Interestingly, we observe that both InfoNCE and MI-InfoNCE achieve identical scores when testing on HMDB51, while MI-InfoNCE outperforms the others on UCF101. This is due to action recognition on UCF101 relies more on static context information.
Selection of Video Generation Target Since our video generating branch is trained with two different objective functions, we present the performance of each separately on the downstream tasks. Therefore, we train models in three different setups, i.e., jointly train CSVR with both $\mathcal{L}_{r}$ and $\mathcal{L}_{p}$, and train CSVR with $\mathcal{L}_{r}$ or $\mathcal{L}_{p}$ independently. As shown in Table 6, the results obtained from the full branch objective surpass the result of reconstruction objective in both cases. Moreover, when compared to the result of prediction objective, the full branch objective outperforms better on the UCF101 dataset and performs equally well on HMDB51.
Hyperparameter investigation of $\sigma_{p},\sigma_{c}$. CSVR contains three branches which are jointly optimized by loss function $L$. $\sigma_{p},\sigma_{c}$ in $L$ are the hyperparameters that coordinate the weights of pose prediction branch and context matching branch, respectively. We perform a grid search on all combinations of two parameters within a range of 0.5 - 1.5 and conduct the pretraining experiments with each combination to find the optimal parameters. Figure 6 shows the results on UCF101 when transferring the pretrained features to the downstream action recognition task. As seen, the best results are achieved when $\sigma_{p}=\sigma_{c}=0.75$, which are chosen for all other experiments.
Table 5: Performance comparison of using different objective functions in video generating branch to train CSVR.
Objective UCF101 HMDB51
$\mathcal{L}_{r}$ 89.2 54.6
$\mathcal{L}_{p}$ 89.9 56.6
$\mathcal{L}_{r}$ $\mathcal{L}_{p}$ 90.3 56.5
Table 6: Action recognition results on UCF101 with different combinations of $\sigma_{p},\sigma_{c}$.
Per-Class Accuracy of Action Recognition with Different Features. To further investigate the contribution of features from different branches to action understanding, we present a per-class accuracy visualization of typical 26 classes from UCF101 in Figure 3. The action classes are divided into three groups shown on horizontal axis. The group represented by orange characters (denoted as ”orange group”) are highly related to dynamic motion. The group represented by blue characters (denoted as ”blue group”) can be easily recognized by static context. And the group represented by black characters (denoted as ”black group”) need both motion and context to distinguish. For the orange group, the dynamic motion feature $f_{d}$ achieves much better results than the static context feature $f_{s}$, while the opposite conclusion is obtained for the blue group. This result demonstrates that the carefully designed pretext tasks can well capture different types of features beneficial to action recognition task. Moreover, by jointly optimizing three branches, the integrated feature $f_{i}$ is refined and greatly improves the feature representation for actions, whether based on motion or context.
Figure 3: Per-class action recognition accuracy with different features. Labels on the horizontal axis are divided into three groups: the orange group are highly related to motion information, the blue group can be easily recognized by static appearance, the black group exhibit both fixed action patterns and rich scene changes. The figure reveals the contributions of different features in action recognition.
4.4 Comparison with State-of-the-art Methods
To verify the effectivenss of our CSVR framework, we conduct experiments on two downstream tasks, i.e., action recognition and video retrieval, with four 3D-CNN backbones, i.e., R(2+1)D, R3D, C3D and S3D. The whole model is pretrained on two datasets, i.e., UCF101 and large-scale Kinetics400, respectively.
Table 7: Action Recognition. Comparison of CSVR with state-of-the-art self-supervised methods on the action recognition task. *STCNet is a specially-designed R3D network used in DynamoNet[162].
Method Backbone Pretrain Dataset Epochs Resolution Clip Length UCF101 HMDB51
DSM[29] C3D UCF101 200 $112\times 112$ 16 70.3 40.5
CoCLR[155] S3D UCF101 700 $128\times 128$ 32 81.4 52.1
MSCL[1] S3D UCF101 200 $128\times 128$ 32 83.7 53.8
DCM[160] R(2+1)D UCF101 120 $112\times 112$ 16 79.7 48.6
VCL[3] R(2+1)D UCF101 200 $112\times 112$ 16 82.1 49.7
TCLR[69] R3D UCF101 400 $112\times 112$ 16 82.4 52.9
MSCL[1] R3D UCF101 600 $112\times 112$ 16 86.7 58.9
[…]
CSVR(Ours) C3D UCF101 120 $112\times 112$ 16 88.2 52.2
CSVR(Ours) S3D UCF101 120 $112\times 112$ 16 88.0 54.8
CSVR(Ours) R(2+1)D UCF101 120 $112\times 112$ 16 90.3 56.5
CSVR(Ours) R3D UCF101 200 $112\times 112$ 16 89.6 60.7
Spation-Temp[9] C3D Kinetics400 20 $112\times 112$ 32 61.2 33.4
DCM[160] C3D Kinetics400 120 $112\times 112$ 16 83.4 52.9
CoCLR[155] S3D Kinetics400 500 $128\times 128$ 32 87.9 54.6
VideoMoCo[14] R(2+1)D Kinetics400 200 $112\times 112$ 32 78.7 49.2
DCM[160] R(2+1)D Kinetics400 120 $112\times 112$ 16 85.7 54.0
FAME[153] R(2+1)D Kinetics400 200 $112\times 112$ 16 84.8 53.5
VCL[3] R(2+1)D Kinetics400 100 $112\times 112$ 16 86.1 54.8
TCLR[69] R3D Kinetics400 100 $112\times 112$ 16 84.1 53.6
TE[18] R3D Kinetics400 200 $128\times 128$ 16 87.1 63.6
DynamoNet[162] STCNet* Kinetics400 200 $112\times 112$ 16 88.1 59.9
[…]
MSCL[1] R3D Kinetics400 400 $112\times 112$ 16 90.7 62.3
CSVR(Ours) C3D Kinetics400 120 $112\times 112$ 16 89.8 60.8
CSVR(Ours) S3D Kinetics400 120 $112\times 112$ 16 88.2 59.2
CSVR(Ours) R(2+1)D Kinetics400 120 $112\times 112$ 16 90.8 63.0
CSVR(Ours) R3D Kinetics400 200 $112\times 112$ 16 92.5 64.5
Table 8: Video Retrieval. Comparison of CSVR with the state-of-the-art on nearest-neighbour video retrieval on UCF101 and HMDB51. Given query test clips, our goal is to find training clips that are from the same class using Recall at $k$(R@$k$) metric.
Method Backbone Epochs UCF101 HMDB51
R@1 R@5 R@10 R@20 R@1 R@5 R@10 R@20
TCGL[16] R(2+1)D 300 21.5 39.3 49.3 59.5 10.5 27.6 39.7 55.6
MCN[158] R(2+1)D 500 52.5 69.5 77.9 83.1 23.7 46.5 58.9 72.4
MotionFit[2] R(2+1)D 120 61.6 75.6 - 85.5 29.4 46.5 - 66.7
DualVar[7] R3D 200 46.7 63.1 69.7 78.0 - - - -
MSCL[1] R3D 400 63.7 79.1 84.0 - 32.6 58.5 70.5 -
CSVR(Ours) R(2+1)D 120 64.4 80.2 87.8 93.1 34.1 57.2 71.0 81.4
CSVR(Ours) R3D 200 64.2 79.5 88.2 92.9 34.5 59.0 71.8 82.2
Action Recognition. We firstly compare CSVR with the previous self-supervised video representation learning methods in terms of the downstream action recognition task. As shown in Table 7, when pretraining on UCF101, our CSVR consistently outperforms all existing methods with the same backbone. Moreover, when pretraining on large-scale dataset Kinetics400, our method also achieves the best results of $92.5\%$ and $64.5\%$ on UCF101 and HMDB51, respectively. We can also observe that the trained model by our CSVR surpasses the methods which use other modality of data like optical flow [155], multi-modal information [160]. Besides, our CSVR needs fewer pretraining epochs than most existing methods. All these results demostrate that CSVR is both effective and efficient for self-supervised video representation learning on action recognition.
Video Retrieval. Besides action recognition, we also conduct evaluations for the downstream video retrieval task on two human action datasets UCF101 and HMDB51. Table 8 shows the results of video retrieval of each method in terms of R@K, where R@K denotes the recall rate of top-K. All methods are pretrained on Kinetics400 with the resolution of 112$\times$112 for fair comparisons. The similar conclusion can be obtained that our CSVR achieves the best results on both UCF101 and HMDB51, showcasing its superiority for self-supervised features learning for human action understanding.
5 Conclusion
This paper designs a self-supervised video representation learning framework named CSVR for action recognition by collaboratively training three branches: generative pose prediction branch, discriminative context matching branch and video generating branch. The pose prediction branch is responsible for extracting dynamic motion features, while the context matching branch focuses on learning static context features. Furthermore, the video generating branch jointly optimizes the dynamic motion and static context features, leading to a more comprehensive feature representation for the downstream tasks. Extensive experiments demonstrate that our CSVR outperforms the state-of-the-art methods on both action recognition and video retrieval tasks. In the future, we will introduce more powerful backbone like Transformer into CSVR to further explore potential of the learning framework for various video understanding tasks. We hope our CSVR can inspire the researches on the collaboration of discriminative and generative self-supervised learning.

### v2

@@
-
+Collaboratively Self-supervised Video Representation Learning for Action Recognition
 Abstract
-Considering the close connection between action recognition and human pose estimation, we design a Collaboratively Self-supervised Video Representation (CSVR) learning framework specific to action recognition by jointly considering generative pose prediction and discriminative context matching as pretext tasks.
+Considering the close connection between action recognition and human pose estimation, we design a Collaboratively Self-supervised Video Representation (CSVR) learning framework specific to action recognition by jointly factoring in generative pose prediction and discriminative context matching as pretext tasks.
 Specifically, our CSVR consists of three branches: a generative pose prediction branch, a discriminative context matching branch, and a video generating branch.
-Among them, the first one encodes dynamic motion feature by utilizing Conditional-GAN to predict the human poses of future frames, and the second branch extracts static context features by pulling the representations of clips and compressed key frames from the same video together while pushing apart the pairs from different videos.
-The third branch is designed to recover the current video frames and predict the future ones, for the purpose of collaboratively improving dynamic motion features and static context features.
-Extensive experiments demonstrate that our method achieves state-of-the-art performance on the UCF101 and HMDB51 datasets.
-keywords:
-Video representation learning, Self-supervised learning, Action recognition, Human pose prediction
-1 Introduction
+Among them, the first one encodes dynamic motion feature by utilizing Conditional-GAN to predict the human poses of future frames, and the second branch extracts static context features by contrasting positive and negative video feature and I-frame feature pairs.
+The third branch is designed to generate both current and future video frames, for the purpose of collaboratively improving dynamic motion features and static context features.
+Extensive experiments demonstrate that our method achieves state-of-the-art performance on multiple popular video datasets.
+Index Terms:
+Video representation learning, Self-supervised learning, Action recognition, Human pose prediction.
+00footnotetext: Manuscript received 5 July 2024; revised 8 December 2024 and 1 January 2025; accepted 2 January 2025.
+Date of publication 16 January 2025; date of current version 15 January 2025.
+This work is partially supported by Natural Science Foundation of China (No.
+U2336213, 62176251) , Strategic Priority Research Program of the Chinese Academy of Sciences (No.
+XDB0680202), Beijing Nova Program (20230484368), Suzhou Frontier Technology Research Project (No.
+SYG202325), and Youth Innovation Promotion Association CAS.
+The associate editor coordinating to the review of this manuscript amd approving it for publication was Prof.
+Zhen Lei.
+(Corresponding author: Shiguang Shan.) Jie Zhang, Zhifan Wan, Lanqing Hu and Shiguang Shan are with the Key Laboratory of AI Safety of CAS, Institute of Computing Technology (ICT), Chinese Academy of Sciences (CAS), Beijing 100190, China.
+They are also affiliated with the School of Computer Science and Technology, University of Chinese Academy of Sciences (UCAS), Beijing 100049, China (e-mail: zhangjie@ict.ac.cn; wanzhifan16@mails.ucas.ac.cn; lanqing.hu@vipl.ict.ac.cn; sgshan@ict.ac.cn).
+Stephen Lin is with Microsoft Research Asia, Beijing 100080, China (e-mail: stevelin@microsoft.com).
+Shuzhe Wu is with Beijing Huawei Digital Technologies Co., Ltd., No.3 Xinxi Road, Haidian District, Beijing 100095, China (e-mail: wushuzhe2@huawei.com).
+I Introduction
 Human action recognition is a fundamental task in computer vision and has gained increasing attention in recent years.
-With a large number of labeled videos, the convolutional neural network (CNN) has achieved great success in human action recognition by learning powerful spatiotemporal representation features [32, 33, 81, 80].
+With a large number of labeled videos, the convolutional neural network (CNN) has achieved great success in human action recognition by learning powerful spatio-temporal representation features [32, 33, 81, 80].
 However, the cost of collecting and tagging large amounts of videos is enormous.
-Fortunately, there is a huge amount of unlabeled videos on the Internet.
+Fortunately, there are a large amount of unlabeled videos on the Internet.
 A common practice is to use a large amount of unsupervised video for robust feature representation learning and then transfer to downstream tasks such as human action recognition, which deserves further exploration.
@@
 This approach has achieved remarkable success in areas such as image recognition, object detection, and image segmentation.
-Besides the self-supervised learning on images [55, 60], there has been growing interest in exploring self-supervised learning for video representation with the increasing availability of large-scale unlabeled video data.
-Early video self-supervised learning approaches mainly focus on effectively exploiting contextual relationships within videos [58, 56].
-Over the years, self-supervised methods have become more powerful and diverse can generally be categorized into two groups: discriminative classification / regression methods [94, 86] and generative prediction ones [29, 6, 4].
+In addition to self-supervised learning on images [55, 60], there has been a growing interest in exploring self-supervised learning for video representation with the increasing availability of large-scale unlabeled video data.
+Early self-supervised video learning approaches focus mainly on effectively exploiting contextual relationships within videos [58, 56, 12].
+Over the years, self-supervised methods have become more powerful and can generally be categorized into two groups: discriminative classification / regression methods [94, 86] and generative prediction methods [168, 29, 6].
 Discriminative methods focus on learning to distinguish between different samples, while generative methods aim to predict missing information.
 Recent efforts have been made to learn general video representations that can be applied to a range of downstream tasks.
-This includes video contrastive learning [94, 18, 69, 1], as well as cross-modal learning [115, 105], which learns from different modalities such as video, audio and text.
+This includes video contrastive learning [94, 18, 69, 1, 65], as well as cross-modal learning [114, 105], which learns from different modalities such as video, audio, and text.
 Although these methods achieve promising results for various video based downstream tasks, they do not explicitly consider designing pretext tasks for action understanding, which may lead to sub-optimal solutions for action recognition.
+Considering the close connection between action recognition and human pose estimation, we design a self-supervised video representation learning framework specific to action recognition by explicitly inducing generative pose prediction as the pretext task.
 One interesting work, DCM [160] designs a two-stream network to separately learn action-related features, i.e., context and motion features, which achieves impressive results for action recognition.
 However, DCM fails to jointly consider the learning of context and motion features, lacking comprehensive view of these two features.
+Differently, we conduct collaboratively self-supervised video feature learning by jointly considering generative pose prediction and discriminative context matching as pretext tasks.
 In this work, we propose the Collaboratively Self-supervised Video Representation (CSVR) framework, which jointly learns dynamic motion and static context features for the downstream action recognition task.
-As illustrated in Figure 1, CSVR consists of three main representation learning branches, i.e., pose prediction branch, context matching branch, and a collaborative training branch named video generating branch.
+As illustrated in Figure 1, CSVR consists of three main representation learning branches, i.e.pose prediction branch, context matching branch, and a collaborative training branch named video generating branch.
 To enhance the performance of action recognition, we resort to the strong relationship between human pose and actions.
@@
 By excluding the potential motion noise from backgrounds, our method can effectively learn dynamic motion features highly related to action.
-Video key frames, known as the intra frame (I-frame) in video compression domain, contain valuable context information.
+Intra frame (I-frame) refers to a type of frame in video compression that contains complete image data and can be used as a reference point for decoding other frames.
+It is typically considered to to contain valuable context information.
 To leverage the abundant pattern in I-frame for action recognition, we incorporate a discriminative context matching branch, which works by pulling the features of clips and I-frames from the same videos closer together, while simultaneously pushing apart pairs from different videos.
@@
 The video generating branch jointly optimizes all the branches, leading to learn more comprehensive video representation.
-Extensive experiments on three popular datasets with various backbones show that our CSVR achieves state-of-the-art performance for two downstream action understanding tasks, i.e., action recognition and video retrieval.
+Extensive experiments on four popular datasets with various backbones show that our CSVR achieves state-of-the-art performance for two downstream action understanding tasks, i.e.action recognition and video retrieval.
 We summarize the contributions in the following:
-1) We introduce a novel framework focusing on action recognition named Collaboratively Self-supervised Video Representation learning (CSVR), which simultaneously learns dynamic motion and static context features using three branches with discriminative and generative pretext tasks.
-2) Considering the strong connections between human pose and actions, we design a generative pose prediction branch that predicts future poses from current pose sequences, enabling efficient and effective extraction of dynamic motion features.
-3) Our carefully designed collaborative video generating branch enables joint optimization of dynamic motion and static context features, resulting in more comprehensive representations that benefit downstream action understanding tasks.
-2 Related Work
+1) Considering the strong connections between human pose and actions, we introduce a novel framework focusing on action recognition named Collaboratively Self-supervised Video Representation learning (CSVR), which simultaneously learns dynamic motion and static context features using three branches with discriminative and generative pretext tasks.
+2) Our carefully designed collaborative video generating branch enables joint optimization of dynamic motion and static context features, resulting in more comprehensive representations that benefit downstream action understanding tasks.
+II Related Work
 As our method is primarily focused on self-supervised video representation learning for action recognition, in this section, we provide a brief overview of some typical works in the fields of video action recognition and self-supervised video representation learning.
-Additionally, we review some works on Generative Adversarial Networks (GAN) [84] as GAN-based generative branches play a crucial role in our method.
-2.1 Video-Based Action Recognition
+Additionally, we review some generation methods as GAN-based [84] generative branches play a crucial role in our method.
+II-A Video-Based Action Recognition
 Videos contain motion information with a complex temporal structure, which requires spatio-temporal feature learning.
@@
 Some of them [43, 44] use pose information as an attention or supervision signal to mine video representations more suitable for action understanding tasks.
-2.2 Self-Supervised Video Representation Learning
+II-B Self-Supervised Video Representation Learning
 Self-supervised video representation learning methods aims to leverage unlabeled video data to learn representations that can be easily transferred to downstream tasks.
@@
 Differently, DCM [160] and P-HLVC [19] aims to learn generalizable features specific to action understanding.
-They simply extract motion features by contrastive learning and leverage diverse motion information, i.e., pose and motion vectors as self-supervised signals.
+They simply extract motion features by contrastive learning and leverage diverse motion information, i.e.pose and motion vectors as self-supervised signals.
 Inspired by these approaches, we propose a Collaboratively Self-supervised Video Representation (CSVR) learning framework that simultaneously considers pose prediction and context matching as pretext tasks.
 By jointly optimizing the whole framework by video generating branch, CSVR provides a more comprehensive feature representation for downstream action recognition task.
-2.3 Generative Adversarial Networks
-Generative adversarial networks (GAN) [84] usually consist of two encoder-decoder architecture networks: a generator producing fake data and a discriminator distinguishing between fake and real data.
+II-C Video Generation
+Over the past years, many efforts are devoted to video generation, which can be generally categorized into three groups: Generative Adversarial Networks (GAN) [84], Variational Autoencoders (VAE) [170] and Diffusion Models [171].
+Generative adversarial network (GAN) [84] is initially used to generate high-fidelity samples that matched the data distribution.
+Some work[14, 168] also employed it as a pretext task for video representation learning.
+GANs usually consist of two encoder-decoder architecture networks: a generator producing fake data and a discriminator distinguishing between fake and real data.
 However, the inherent randomness of GAN output makes it difficult to control the generated data.
@@
 TGAN v2[77] is a video generation network that can efficiently generate realistic videos.
-Some works [163, 164] prove the feasibility of learning video representation from generative learning methods.
-We believe the process of generating diverse samples utilizing GANs can better capture human motion and learn representations which benefit the action understanding tasks.
-In CSVR, we employ 1D-CNN-based CGAN in pose prediction branch and R3D-based TGAN v2 in video generating branch.
+Some works [14, 163, 164, 168] prove the feasibility of learning video representation from generative learning methods.
+VideoMoCo[14] is the first work to use GANs for video representation learning.
+It reconstructs the foreground and background of videos separately using two GANs and applies the encoder of generator to action recognition downstream tasks.
+TS-GAN[168] employs a spatio-temporal two-stream GAN network to reconstruct keyframes and optical flow information, leveraging the learned features in downstream tasks.
+Unfortunately, both approaches underperformed across multiple datasets.
+Although these methods using GANs can effectively reconstruct video dynamics, their features do not perform well in downstream classification tasks.
+Variational Autoencoders (VAEs) [170] have become widely adopted in two-stage generative models, where the first stage compresses images into a lower-dimensional latent representation, and the second stage generates images from this latent space.
+In the context of video generation, 2D VAEs are often extended to 3D variants by incorporating 3D convolutional layers or temporal attention mechanisms [172, 173, 174, 185].
+These approaches have demonstrated promising results in video generation.
+More recently, diffusion models have achieved remarkable success in both image [175, 176, 177, 178] and video generation [179, 180, 181, 182, 183, 184, 186].
+Video diffusion models can be broadly categorized into pixel-space methods [179, 183] and latent-space methods [181, 182].
+Compared to the pixel-space methods, latent-space methods offer significant efficiency advantages for video generation.
+However, they still incur high computational costs, which makes them challenging to apply in self-supervised video representation learning.
+Therefore, in our CSVR, we employ 1D-CNN-based CGAN in pose prediction branch and R3D-based TGAN v2 in video generating branch.
 To enhance the stability and efficiency of our training process, we also adopt the gradient penalty from WGAN-GP [72] and the parameter update strategy from OSGANs [50].
 These carefully-designed objectives help our method efficiently generate more accurate and diverse samples, thus improve the quality of learned dynamic motion representations.
-3 Methodology
-3.1 Overall Framework
+III Methodology
+III-A Overall Framework
 In this paper, we propose a novel framework for action recognition called Collaboratively Self-Supervised Video Representation learning (CSVR).
@@
 We will introduce the detail of each branch in the following sections.
-3.2 Generative Pose Prediction Branch
-The generative pose prediction branch is conducted to learn dynamic motion features, which consists of two modules, i.e., a pose generator ($G_{p}$) and a pose discriminator ($D_{p}$), similar to the structure of CGAN [83].
+III-B Generative Pose Prediction Branch
+The generative pose prediction branch is conducted to learn dynamic motion features, which consists of two modules, i.e.a pose generator ($G_{p}$) and a pose discriminator ($D_{p}$), similar to the structure of CGAN [83].
 Using the poses from current video clips as input, the objective of pose generator ($G_{p}$) is to generate plausible future poses.
@@
 Suppose $p_{t}\in\mathbb{R}^{2N}$ denotes the human pose captured in the $t$-th video frame, where $N$ is the number of skeleton joints.
-Given a randomly selected $T$-frame video clip, we apply OpenPose human skeleton detector [79] to estimate the current pose sequence $P_{in}=(p_{0},p_{1},\cdots,p_{T-1})\in\mathbb{R}^{2N}$ and the future pose sequence $P_{gt}=(p_{T},p_{T+1},...,p_{T+T^{\prime}-1})$ of length $T^{\prime}$ frames.
+Given a randomly selected $T$-frame video clip, we apply OpenPose human skeleton detector [79] to estimate the current pose sequence $P_{in}=(p_{0},p_{1},\cdots,p_{T-1})\in\mathbb{R}^{1\times(T\times 2N)}$ and the future pose sequence $P_{gt}=(p_{T},p_{T+1},...,p_{T+T^{\prime}-1})$ of length $T^{\prime}$ frames.
 The input to $G_{p}$ is a latent noise vector $z_{p}$ and the current pose sequence $P_{in}$.
@@
 The objective function is formulated as:
-$\begin{split}\min_{G_{p}}&\max_{D_{p}}\mathcal{L}_{pose}=\\ &E_{P_{gt}}[\log D_{p}(P_{gt})]+E_{z}[\log(1-D_{p}(G_{p}(z_{p}|P_{in}))]+\\ &\frac{1}{T^{\prime}}\sum^{T+T^{\prime}-1}_{t=T}||p_{t}-p^{\prime}_{t}||_{2}^{2}+E_{p}[(\parallel\bigtriangledown D_{p}(P|P_{in})\parallel_{2}-1)^{2}]\end{split}$ (1)
+$\begin{split}\min_{G_{p}}&\max_{D_{p}}\mathcal{L}_{pose}=\\ &\mathbb{E}_{p\sim P_{gt}}[\log D_{p}(P_{gt})]+\mathbb{E}_{z\sim p(z)}[\log(1-D_{p}(G_{p}(z_{p}|P_{in}))]+\\ &\frac{1}{T^{\prime}}\sum^{T+T^{\prime}-1}_{t=T}||p_{t}-p^{\prime}_{t}||_{2}^{2}+\mathbb{E}_{p}[(\parallel\bigtriangledown D_{p}(P|P_{in})\parallel_{2}-1)^{2}]\end{split}$ (1)
 The objective function mainly consists of two parts: L2 reconstruction loss contributes to the generation of accurate future poses, and the adversarial loss helps $G_{p}$ to learn the underlying distribution of the input data, leading to more realistic poses.
@@
 Subsequently, we add convolutional layers and MLPs after the last layer in encoder of $G_{p}$ to keep the dimension of dynamic motion feature $f_{d}$ align with static context feature $f_{s}$.
-3.3 Discriminative Context Matching Branch
+III-C Discriminative Context Matching Branch
 Besides the dynamic motion information, the context information is also significant for human action analysis.
@@
 We extract features $x_{i}\in\mathbb{R}^{C_{1}\times T_{1}\times H_{1}\times W_{1}}$ from a random video clip $i$ and feature $z_{i}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ from the corresponding I-frame of the same video, where $C_{1}$, $T_{1}$, $H_{1}$, and $W_{1}$ denote the channel, time, height, and width of video clips, respectively.
-We use global average pooling to obtain their global representations $x_{i}^{pool}\in\mathbb{R}^{C_{1}}$ and $z_{i}^{pool}\in\mathbb{R}^{C_{2}}$ We then apply two-layer MLP heads, $g^{V}$ and $g^{I}$, by taking the features of the video clip and I-frame as inputs, i.e., $x_{i}^{*}=g^{V}(x_{i}^{pool})\in\mathbb{R}^{C}$ and $z_{i}^{*}=g^{I}(z_{i}^{pool})\in\mathbb{R}^{C}$.
+We use global average pooling to obtain their global representations $x_{i}^{pool}\in\mathbb{R}^{C_{1}}$ and $z_{i}^{pool}\in\mathbb{R}^{C_{2}}$ We then apply two-layer MLP heads, $g^{V}$ and $g^{I}$, by taking the features of the video clip and I-frame as inputs, i.e.$x_{i}^{*}=g^{V}(x_{i}^{pool})\in\mathbb{R}^{C}$ and $z_{i}^{*}=g^{I}(z_{i}^{pool})\in\mathbb{R}^{C}$.
 Finally, the Multi-Instance-InfoNCE (MI-InfoNCE) loss [161] is applied to pull the video clip and I-frame feature pairs from the same video closer and push pairs from different video further away:
-$\begin{split}\mathcal{L}_{context}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\sum\exp(\frac{\cos(z^{*}_{i},x^{*}_{i})}{\tau})}{\sum\exp(\frac{\cos(z^{*}_{i},x^{*}_{i})}{\tau})+\sum_{k\neq i}^{B}\exp(\frac{\cos(z^{*}_{k},x^{*}_{i})}{\tau})}\end{split}$ (2)
+$\begin{split}&\mathcal{L}_{context}=\\ &-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\sum\exp(\frac{\cos(z^{*}_{i},x^{*}_{i})}{\tau})}{\sum\exp(\frac{\cos(z^{*}_{i},x^{*}_{i})}{\tau})+\sum_{k\neq i}^{B}\exp(\frac{\cos(z^{*}_{k},x^{*}_{i})}{\tau})}\end{split}$ (2)
 where $z_{i}^{*}$ and $z_{k}^{*}$ denote the positive and negative I-frame features, $B$ denotes the number of samples in the minibatch, $\cos(z,x)=(z^{T}x)/(||z||_{2}\cdot||x||_{2})$ denotes the cosine similarity between $z$ and $x$, and $\tau$ is a temperature adjusting the scale of cosine similarities.
@@
 After pretraining on large video dataset, the video feature $x_{i}^{*}$ from the last layer of V-Network will serve as the static context feature $f_{s}\in\mathbb{R}^{C}$ in our model.
-3.4 Collaborative Video Generating Branch
+III-D Collaborative Video Generating Branch
 After achieving the dynamic motion feature $f_{d}$ and tatic context feature $f_{s}$, we further conduct a collaborative video generating branch based on TGAN v2 [77] to learn more comprehensive video features by fusing these two features.
@@
 Inspired by the prior work [119, 118], we adopt Adaptive Instance Normalization (AdaIN) for feature fusion.
-We firstly align the dimensions of the dynamic motion feature $f_{d}$ and static context feature $f_{s}$ by MLPs.
-The fusion of $f_{d}$ and $f_{s}$ is accomplished through a denormalization process.
+We firstly align the dimensions of the dynamic motion feature $f_{d}$ from the pose prediction branch and static context feature $f_{s}$ from V-network by adding convolutional layers and MLPs after the encoder of $G_{p}$.
+Then, the fusion of $f_{d}$ and $f_{s}$ is accomplished through a denormalization process.
 The detailed structure of the feature integration layer is shown in Fig.
@@
 The loss objective for video generating branch $\mathcal{L}_{video}$ involve two objectives $\mathcal{L}_{r}$ and $\mathcal{L}_{p}$:
-$\begin{split}\min_{G^{r}_{v}}&\max_{D^{r}_{v}}\mathcal{L}_{r}=\\ &E_{V_{r}}[\log D^{r}_{v}(V_{r})]+E_{z}[\log(1-D^{r}_{V}(G^{r}_{V}(z^{r}_{v}|f_{i}))]+\\ &\frac{1}{THW}\sum^{THW}||v_{t}^{r}(h,w)-g_{t}^{r}(h,w)||_{2}^{2}\end{split}$ (6)
-$\begin{split}\min_{G^{p}_{v}}&\max_{D^{p}_{v}}\mathcal{L}_{p}=\\ &E_{V_{p}}[\log D^{p}_{v}(V_{p})]+E_{z}[\log(1-D^{p}_{V}(G^{p}_{V}(z_{v}^{p}|f_{i}))]+\\ &\frac{1}{THW}\sum^{THW}||v_{t}^{p}(h,w)-g_{t}^{p}(h,w)||_{2}^{2}\end{split}$ (7)
+$\begin{split}\min_{G^{r}_{v}}&\max_{D^{r}_{v}}\mathcal{L}_{r}=\\ &\mathbb{E}_{v\sim V_{r}}[\log D^{r}_{v}(V_{r})]+\mathbb{E}_{z\sim p(z)}[\log(1-D^{r}_{V}(G^{r}_{V}(z^{r}_{v}|f_{i}))]+\\ &\frac{1}{THW}\sum^{THW}||v_{t}^{r}(h,w)-g_{t}^{r}(h,w)||_{2}^{2}\end{split}$ (6)
+$\begin{split}\min_{G^{p}_{v}}&\max_{D^{p}_{v}}\mathcal{L}_{p}=\\ &\mathbb{E}_{v\sim V_{p}}[\log D^{p}_{v}(V_{p})]+\mathbb{E}_{z\sim p(z)}[\log(1-D^{p}_{V}(G^{p}_{V}(z_{v}^{p}|f_{i}))]+\\ &\frac{1}{THW}\sum^{THW}||v_{t}^{p}(h,w)-g_{t}^{p}(h,w)||_{2}^{2}\end{split}$ (7)
 $\begin{split}\mathcal{L}_{video}=(\mathcal{L}_{r}+\mathcal{L}_{p})/2\end{split}$ (8)
 where $V_{r}$ and $V_{p}$ denote the generation target of each generator videos and $T$, $H$, $W$ represent the temporal and spatial dimensions of video clip, respectively.
-$v_{t}^{r}(h,w)$ and $v_{t}^{p}(h,w)$ indicates the pixel value at the $(h,w)$ coordinates on the $t$-th frame in the ground truth video.
+$v_{t}^{r}(h,w)$ and $v_{t}^{p}(h,w)$ indicate the pixel value at the $(h,w)$ coordinates on the $t$-th frame in the ground truth video.
 The generated videos are denoted as $g_{t}^{r}(h,w)$ and $g_{t}^{p}(h,w)$.
@@
 Two parallel training objectives enhance the understanding of the temporal relationship of the video, leading to a more comprehensive video representation.
-3.5 Joint Optimization
+III-E Joint Optimization
 In our CSVR, we jointly optimize the three branches by linearly combining the losses from each branch:
@@
 where the $\sigma_{p},\sigma_{c}$ are scalar hyper-parameters.
-All the networks in three branches are jointly optimized with $\mathcal{L}$ in an end-to-end fashion.
-Finally, for fine-tuning on downstream tasks, we utilize the pretrained parameters from the V-Network as the initialization.
-4 Experiments
-In this section, we conduct experiments on three popular action recognition datasets, i.e., UCF101 [121], HMDB51 [122], and Kinetics400[148] to verify the effectiveness of our CSVR.
+Since our CSVR contains three branches, we firstly pretrain each branch separately and then jointly optimize them with $\mathcal{L}$ in an end-to-end fashion for stable training.
+Finally, for fine-tuning on downstream tasks, we utilize the pretrained integrated feature $f_{i}$ as the initialization.
+IV Experiments
+In this section, we conduct experiments on four popular action recognition datasets, i.e.UCF101 [120], HMDB51 [121], Something-Something v2(SSv2) [169] and Kinetics-400 [148] to verify the effectiveness of our CSVR.
 Besides, we also conduct ablation studies to investigate the significance of each branch in our method.
-4.1 Dataset
-The experiments of CSVR are mainly conducted on three datasets.
+IV-A Dataset
+The experiments of CSVR are mainly conducted on four datasets.
 UCF101 contains 13,320 videos covering 101 action classes, while HMDB51 consists of 6,766 videos in 51 classes.
 Following previous methods [94, 160], we use split 1 of UCF101 and HMDB51 in our experiments.
-Kinetics400 is a large-scale dataset containing 246K videos in its training set which includes 400 classes of human actions.
-During the experiments, we pretrain our CSVR on UCF101 training set and Kinetics400 training set, respectively.
-Then we finetune the pretrained V-Network to two downstream tasks, i.e., action recognition and video retrieval, on UCF101 and HMDB51, respectively.
-4.2 Implementation Details
+SSv2 is another extensive video dataset, emphasizes temporal information with its 169K training and 20K validation videos across 174 action classes.
+Kinetics-400 is a large-scale dataset containing 246K videos in its training set which includes 400 classes of human actions.
+During the experiments, we pretrain our CSVR on UCF101 training set and Kinetics-400 training set, respectively.
+Then we finetune the network initialized with the pretrained $f_{i}$ to two downstream tasks, i.e., action recognition and video retrieval, on UCF101 and HMDB51, respectively.
+IV-B Implementation Details
 Backbone For context matching branch, we explore four 3D-CNN backbones as V-Network, i.e., C3D [149], R(2+1)D[150], R3D [151, 80], and S3D [152] to extract video representations.
@@
 In addition, R3D-based TGAN v2 is leveraged in video generating branch.
-It should be noted that during inference, the major computation cost is from video representation extraction V-Network which is mainly determined by various 3D-CNN backbones.
+It should be noted that during inference, the major computation cost is from video representation extraction network which is mainly determined by various 3D-CNN backbones.
 Pretraining Since our CSVR consists of three branches, we firstly pretrain each branch separately and then jointly optimize them in self-supervised manner for stable training.
@@
 As for the generation of $f_{i}$, the feature of the last layer in pose generator is utilized as the the dynamic motion feature $f_{d}$.
-After investigating the selection of hyper parameters, which is introduced in Section 4.3, we set $\sigma_{p}=\sigma_{c}=0.75$ in our final experiment.
-Finetuning After pretraining our CSVR on video dataset, the integration feature $f_{i}$ from the V-Network is utilized to finetune on downstream tasks.
+After investigating the selection of hyper parameters, which is introduced in Section IV-C, we set $\sigma_{p}=\sigma_{c}=0.75$ in our final experiment.
+Finetuning After pretraining our CSVR on video dataset, the integrated feature $f_{i}$ is utilized to finetune on downstream tasks.
 We conduct experiments on two widely-used human action datasets, i.e., UCF101 and HMDB51, and report Top-1 accuracy of action recognition task and Recall at k (R@k) of video retrieval task on each dataset.
@@
 For HMDB51, we set the learning rate to $lr=0.0002\times B$ and the weight decay to 0.002.
+Pretrain models are finetuned on both datasets for 100 epochs.
 We incorporate a dropout ratio of 0.3 in the second-to-last layer of the network during finetuning.
@@
 For this task, we do not apply any augmentation except for random crop.
-4.3 Ablation Study
+IV-C Ablation Study
 To further investigate the effectiveness of each branch of CSVR, we conduct experiments by firstly pretraining on UCF101 training set and then finetuning on the same set for action recognition.
-We choose R(2+1)D as the backbone of V-network for all ablation study experiments.
-Table 1: Action recognition performance comparison of different branches.
-4-th row represents the integrated feature $f_{i}$ without optimization from video generating branch, while in the last row, $f_{i}$ is learned by the full CSVR pretraining framework.
-$f_{d}$ $f_{s}$ $f_{i}$ $V$ UCF101 HMDB51
+We choose R(2+1)D as the backbone of V-network and finetuning network for all ablation studies.
+TABLE I: Action recognition performance comparison of different branches.
+In table below, $f_{i}$(AdaIN) and $f_{i}$(Concate) represents the integrated feature $f_{i}$ generated by AdaIN or directly concatenating, both of which are generated without optimization from video generating branch.
+$f_{i}$(full) indicates $f_{i}$ learned by the full CSVR pretraining framework.
+$f_{d}$ $f_{s}$ $f_{i}$(AdaIN) $f_{i}$(Concat) $f_{i}$(full) UCF101 HMDB51
 (Training from scratch) 65.0 32.5
@@
 81.0 46.3
-90.3 56.5
-Table 2: Correlation between pretext and downstream tasks.
+75.2 41.7
+90.4 56.6
+Effectiveness of Static and Dynamic Features.
+Since both static information and dynamic motion are crucial to understanding actions, we evaluate the contribution of features from the pose prediction branch and the context matching branch to the performance improvement in this experiment.
+Table I shows the final action recognition performance on UCF101 and HMDB51 with features from various branches as the initialization for finetuning on downstream action recognition.
+As seen, both the static context feature $f_{s}$ and dynamic motion feature $f_{d}$ significantly improve the action recognition accuracy compared to the baseline of training from scratch, demonstrating the superiority of designing context matching and pose prediction branches.
+Interestingly, the dynamic motion feature $f_{d}$ achieves better results than static context feature $f_{s}$ on HMDB51, while $f_{d}$ underperform $f_{s}$ on UCF101.
+This may be because most classes in UCF101 can be classified by context information, while classes in HMDB51 are mainly distinguished by the motion information.
+Effectiveness of Video Generating Branch.
+The collaborative video generating branch contains two components: the integration layer and the video generating networks.
+Following previous practice, we evaluate the effectiveness of these components by finetuning the feature on action recognition tasks on UCF101 and HMDB51.
+By default, the integrated feature $f_{i}$ is generated by fusing the static context feature $f_{s}$ and dynamic motion feature $f_{d}$ with AdaIN.
+As seen in Table.
+I, when employing $f_{i}$ for the downstream task finetuning, the improvement on accuracy further illustrates the complementarity of static context and dynamic motion features for action recognition.
+To further emphasize the effectiveness of AdaIN, we directly concatenate $f_{d}$ and $f_{s}$, followed by convolution to generate the concat feature $f_{i}$(Concat).
+It can be observed that the concat feature offers slight improvements over those produced by previous branches in downstream tasks and performs much worse than the feature generated by AdaIN.
+Finally, we use the full CSVR framework to further improve the integrated feature $f_{i}$ by jointly optimizing three branches.
+The accuracy on UCF101 and HMDB51 is significantly improved to 90.3% and 56.5%, respectively, which shows the effectiveness of collaboratively learning static context and dynamic motion features for action recognition.
+TABLE II: Correlation between pretext and downstream tasks.
 The linear relationship between the results indicates the effectiveness of our pretraining.
@@
 […]
-Effectiveness of Different Branches.
-Since our method includes three different branches, we evaluate the contribution of each branch to the performance improvement.
-Table 1 shows the final action recognition performance on UCF101 and HMDB51 with features from various branches as the initialization for finetuning on downstream action recognition.
-For inference, we transfer different latent features to V-Network and finetune them on downstream tasks.
-As seen, both the static context feature $f_{s}$ and dynamic motion feature $f_{d}$ significantly improve the action recognition accuracy compared to the baseline of training from scratch, which demonstrate the superiority of designing context matching and pose prediction branches.
-Interestingly, the dynamic motion feature $f_{d}$ achieves better results than static context feature $f_{s}$ on HMDB51, while $f_{d}$ underperform $f_{s}$ on UCF101.
-This may be because most classes in UCF101 can be classified by context information, while classes in HMDB51 are mainly distinguished by the motion information.
-Besides, we achieve the integrated feature $f_{i}$ by fusing the static context feature $f_{s}$ and dynamic motion feature $f_{d}$ with AdaIN.
-By employing $f_{i}$ for the downstream task finetuning, the improvement on accuracy further illustrates the complementarity of static context and dynamic motion features for action recognition.
-Finally, we introduce video generating branch to further improve the integrated feature $f_{i}$ by jointly optimizing three branches.
-The accuracy on UCF101 and HMDB51 is significantly improved to 90.3% and 56.5%, respectively, which shows the effectiveness of collaboratively learning static context and dynamic motion features for action recognition.
 Correlation between pretext and downstream tasks.
@@
 The performance measurements include the top-1 accuracy of context matching, mean squared error (MSE) of pose prediction, inception score (IS) of generated videos and top-1 accuracy of downstream action recognition.
-The results in Table 2 show a consistent improvement in action recognition accuracy as the performance of the pretext tasks increases, which indicates the strong correlations between various pretext and downstream tasks.
-Table 3: Comparison of $f_{d}$ from different layers of $G_{p}$ encoder.
-Table 4: Comparison of $f_{s}$ with different loss functions.
+The results in Table II show a consistent improvement in action recognition accuracy as the performance of the pretext tasks increases, which indicates the strong correlations between various pretext and downstream tasks.
+TABLE III: Comparison of feature $f_{d}$ from different layers of $G_{p}$ encoder.
 Layer UCF101 HMDB51
@@
 Layer 6 75.1 38.4
+TABLE IV: Comparison of feature $f_{s}$ with different loss functions.
 Objective UCF101 HMDB51
@@
 MI-InfoNCE 73.1 40.2
-Table 4: Comparison of $f_{s}$ with different loss functions.
 Selection of Dynamic Motion Feature.
 In order to locate the hidden layer which enriches the most dynamic information, we pretrain the pose generator $G_{p}$ on poses from UCF101 dataset and use features from different layers of encoder in $G_{p}$ as dynamic motion features $f_{d}$.
-We report the results of different $f_{d}$s finetuned on action recognition task.
-As shown in Table 4, $f_{d}$ from the last layer achieves the best performance.
+We report the results of different $f_{d}$ finetuned on action recognition task.
+As shown in Table IV, $f_{d}$ from the last layer achieves the best performance.
 Selection of Contrastive Loss.
@@
 The widely used InfoNCE loss [117] compares the feature with limited positive samples and multiple negative samples, which can be formalized as:
-$\begin{split}\mathcal{L}_{context}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\frac{\cos(z^{*}_{i},x_{i})}{\tau})}{\exp(\frac{\cos(z^{*}_{i},x_{i})}{\tau})+\sum_{k\neq i}^{B}\exp(\frac{\cos(z^{*}_{k},x_{i})}{\tau})}\end{split}$ (10)
+$\begin{split}&\mathcal{L}_{context}=\\ &-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\frac{\cos(z^{*}_{i},x_{i})}{\tau})}{\exp(\frac{\cos(z^{*}_{i},x_{i})}{\tau})+\sum_{k\neq i}^{B}\exp(\frac{\cos(z^{*}_{k},x_{i})}{\tau})}\end{split}$ (10)
 where $x_{i}$ denotes the global video representation, and $z_{i}^{*}$ and $z_{k}^{*}$ denote the positive and negative I-frame features, other symbols are the same as Equation 2.
@@
 When calculating InfoNCE and L2 loss, we randomly select an I-frame from the video clip as learning objective to ensure that the loss function produces correct outputs.
-As shown in Table 4, MI-InfoNCE achieve optimal results among three losses.
+As shown in Table IV, MI-InfoNCE achieve optimal results among three losses.
 Interestingly, we observe that both InfoNCE and MI-InfoNCE achieve identical scores when testing on HMDB51, while MI-InfoNCE outperforms the others on UCF101.
@@
 Therefore, we train models in three different setups, i.e., jointly train CSVR with both $\mathcal{L}_{r}$ and $\mathcal{L}_{p}$, and train CSVR with $\mathcal{L}_{r}$ or $\mathcal{L}_{p}$ independently.
-As shown in Table 6, the results obtained from the full branch objective surpass the result of reconstruction objective in both cases.
+As shown in Table V, the results obtained from the full branch objective surpass the result of reconstruction objective in both cases.
 Moreover, when compared to the result of prediction objective, the full branch objective outperforms better on the UCF101 dataset and performs equally well on HMDB51.
+When conducting experiments on the SSv2 dataset, which is more sensitive to temporal information, the performance using the full branch objective significantly surpassed that of using $\mathcal{L}_{r}$ or $\mathcal{L}_{p}$ independently.
+TABLE V: Performance comparison of using different objective functions in video generating branch to train CSVR.
+Objective UCF101 HMDB51 SSv2
+$\mathcal{L}_{r}$ 89.2 54.6 56.1
+$\mathcal{L}_{p}$ 89.9 56.6 55.3
+$\mathcal{L}_{r}$ $\mathcal{L}_{p}$ 90.4 56.6 59.2
+Figure 3: Action recognition results on UCF101 with different combinations of $\sigma_{p},\sigma_{c}$.
 Hyperparameter Investigation of $\sigma_{p},\sigma_{c}$.
@@
 $\sigma_{p},\sigma_{c}$ in $L$ are the hyperparameters that coordinate the weights of pose prediction branch and context matching branch, respectively.
-We perform a grid search on all combinations of two parameters within a range of 0.5 - 1.5 and conduct the pretraining experiments with each combination to find the optimal parameters.
+We perform a grid search on all combinations of two parameters within a range of 0.5 - 1.25 and conduct the pretraining experiments with each combination to find the optimal parameters.
 Figure 3 shows the results on UCF101 when transferring the pretrained features to the downstream action recognition task.
 As seen, the best results are achieved when $\sigma_{p}=\sigma_{c}=0.75$, which are chosen for all other experiments.
-Table 5: Performance comparison of using different objective functions in video generating branch to train CSVR.
-Objective UCF101 HMDB51
-$\mathcal{L}_{r}$ 89.2 54.6
-$\mathcal{L}_{p}$ 89.9 56.6
-$\mathcal{L}_{r}$ $\mathcal{L}_{p}$ 90.3 56.5
-Table 6: Action recognition results on UCF101 with different combinations of $\sigma_{p},\sigma_{c}$.
 Per-Class Accuracy of Action Recognition with Different Features.
@@
 The figure reveals the contributions of different features in action recognition.
-4.4 Comparison with State-of-the-art Methods
+IV-D Comparison with State-of-the-art Methods
 To verify the effectivenss of our CSVR framework, we conduct experiments on two downstream tasks, i.e., action recognition and video retrieval, with four 3D-CNN backbones, i.e., R(2+1)D, R3D, C3D and S3D.
-The whole model is pretrained on two datasets, i.e., UCF101 and large-scale Kinetics400, respectively.
-Table 7: Action Recognition. Comparison of CSVR with state-of-the-art self-supervised methods on the action recognition task. *STCNet is a specially-designed R3D network used in DynamoNet[162].
-Method Backbone Pretrain Dataset Epochs Resolution Clip Length UCF101 HMDB51
+The whole model is pretrained on two datasets, i.e., UCF101 and large-scale Kinetics-400, respectively.
+TABLE VI: Action Recognition on UCF101 and HMDB51.
+Comparison of CSVR with state-of-the-art self-supervised methods on the action recognition task.
+*STCNet is a specially-designed R3D network used in DynamoNet[162].
+{tabu}
+lc—cccc—cc Method Backbone Pretrain Dataset Epochs Resolution Clip Length UCF101 HMDB51
 DSM[29] C3D UCF101 200 $112\times 112$ 16 70.3 40.5
@@
 MSCL[1] R3D UCF101 600 $112\times 112$ 16 86.7 58.9
-[…]
-CSVR(Ours) C3D UCF101 120 $112\times 112$ 16 88.2 52.2
-CSVR(Ours) S3D UCF101 120 $112\times 112$ 16 88.0 54.8
-CSVR(Ours) R(2+1)D UCF101 120 $112\times 112$ 16 90.3 56.5
+Vi2CLR[161] S3D UCF101 300 $112\times 112$ 16 82.8 52.9
+CSVR(Ours) C3D UCF101 200 $112\times 112$ 16 88.3 52.3
+CSVR(Ours) S3D UCF101 200 $112\times 112$ 16 88.1 54.9
+CSVR(Ours) R(2+1)D UCF101 200 $112\times 112$ 16 90.4 56.6
 CSVR(Ours) R3D UCF101 200 $112\times 112$ 16 89.6 60.7
-Spation-Temp[9] C3D Kinetics400 20 $112\times 112$ 32 61.2 33.4
-DCM[160] C3D Kinetics400 120 $112\times 112$ 16 83.4 52.9
-CoCLR[155] S3D Kinetics400 500 $128\times 128$ 32 87.9 54.6
-VideoMoCo[14] R(2+1)D Kinetics400 200 $112\times 112$ 32 78.7 49.2
-DCM[160] R(2+1)D Kinetics400 120 $112\times 112$ 16 85.7 54.0
-FAME[153] R(2+1)D Kinetics400 200 $112\times 112$ 16 84.8 53.5
-VCL[3] R(2+1)D Kinetics400 100 $112\times 112$ 16 86.1 54.8
-TCLR[69] R3D Kinetics400 100 $112\times 112$ 16 84.1 53.6
-TE[18] R3D Kinetics400 200 $128\times 128$ 16 87.1 63.6
-DynamoNet[162] STCNet* Kinetics400 200 $112\times 112$ 16 88.1 59.9
-[…]
-MSCL[1] R3D Kinetics400 400 $112\times 112$ 16 90.7 62.3
-CSVR(Ours) C3D Kinetics400 120 $112\times 112$ 16 89.8 60.8
-CSVR(Ours) S3D Kinetics400 120 $112\times 112$ 16 88.2 59.2
-CSVR(Ours) R(2+1)D Kinetics400 120 $112\times 112$ 16 90.8 63.0
-CSVR(Ours) R3D Kinetics400 200 $112\times 112$ 16 92.5 64.5
-Table 8: Video Retrieval.
+DCM[160] C3D Kinetics-400 120 $112\times 112$ 16 83.4 52.9
+CoCLR[155] S3D Kinetics-400 500 $128\times 128$ 32 87.9 54.6
+VideoMoCo[14] R(2+1)D Kinetics-400 200 $112\times 112$ 32 78.7 49.2
+DCM[160] R(2+1)D Kinetics-400 120 $112\times 112$ 16 85.7 54.0
+FAME[153] R(2+1)D Kinetics-400 200 $112\times 112$ 16 84.8 53.5
+VCL[3] R(2+1)D Kinetics-400 100 $112\times 112$ 16 86.1 54.8
+TCLR[69] R3D Kinetics-400 100 $112\times 112$ 16 84.1 53.6
+TE[18] R3D Kinetics-400 200 $128\times 128$ 16 87.1 63.6
+DynamoNet[162] STCNet* Kinetics-400 200 $112\times 112$ 16 88.1 59.9
+Vi2CLR[161] S3D Kinetics-400 300 $112\times 112$ 16 89.1 55.7
+MSCL[1] R3D Kinetics-400 400 $112\times 112$ 16 90.7 62.3
+KnowledgeGuide[167] C3D Kinetics-400 400 $112\times 112$ 16 82.7 47.0
+KnowledgeGuide[167] R(2+1)D Kinetics-400 400 $112\times 112$ 16 82.4 46.7
+CSVR(Ours) C3D Kinetics-400 200 $112\times 112$ 16 92.8 62.8
+CSVR(Ours) S3D Kinetics-400 200 $112\times 112$ 16 90.2 61.2
+CSVR(Ours) R(2+1)D Kinetics-400 200 $112\times 112$ 16 92.8 64.0
+CSVR(Ours) R3D Kinetics-400 200 $112\times 112$ 16 94.5 65.5
+Action Recognition on UCF101 and HMDB51.
+We firstly compare CSVR with the previous self-supervised video representation learning methods in terms of the downstream action recognition task.
+As shown in Table IV-D, when pretraining on UCF101, our CSVR consistently outperforms all existing methods with the same backbone.
+Moreover, when pretraining on large-scale dataset Kinetics-400, our method also achieves the best results of $94.5\%$ and $65.5\%$ on UCF101 and HMDB51, respectively.
+We can also observe that the trained model by our CSVR surpasses the methods which use other modality of data like optical flow [155], multi-modal information [160].
+Besides, our CSVR needs fewer pretraining epochs than most existing methods.
+All these results demostrate that CSVR is both effective and efficient for self-supervised video representation learning on action recognition.
+TABLE VII: Action Recognition on SSv2. Comparison of CSVR with state-of-the-art self-supervised methods on the action recognition task.
+{tabu}
+l—ccccc—cc Method Backbone Pretrain Dataset Epochs Resolution Clip Length Top-1
+VideoMoCo[14] R(2+1)D Kinetics-400 200 $112\times 112$ 32 19.5
+$\rho$MoCo[65] SlowFast-R50 Kinetics-400 200 $224\times 224$ 16 54.4
+CORP[166] I3D Kinetics-400 800 N/A 16 48.8
+SVT[165] ViT-B Kinetics-400 20 $224\times 224$ 16 59.2
+CSVR(Ours) R(2+1)D Kinetics-400 200 $112\times 112$ 16 59.2
+Action Recognition on SSv2.
+Comparing to UCF101 and HMDB51, SSv2 exhibits less scene bias than previous datasets and require a stronger temporal understanding to accurately classify actions.
+As shown in Table.
+IV-D, our method outperforms the best previous methods using 3D-CNN backbones by an absolute margin of 4.8% on SSv2 when pretrained on Kinetics-400.
+In addition, the performance of CSVR on the SSv2 action recognition task is comparable to that of SVT[165], which uses a stronger ViT backbone[89] and higher video resolution.
+Such results demonstrating its suitability for video datasets captured in controlled, real-world settings.
+TABLE VIII: Video Retrieval.
 Comparison of CSVR with the state-of-the-art on nearest-neighbour video retrieval on UCF101 and HMDB51.
@@
 MSCL[1] R3D 400 63.7 79.1 84.0 - 32.6 58.5 70.5 -
-CSVR(Ours) R(2+1)D 120 64.4 80.2 87.8 93.1 34.1 57.2 71.0 81.4
+KnowledgeGuide[167] R(2+1)D 400 52.6 70.1 77.4 85.2 22.2 45.9 59.9 74.2
+KnowledgeGuide[167] C3D 400 49.9 68.3 77.4 84.9 22.4 46.9 60.3 73.1
+CSVR(Ours) R(2+1)D 200 64.4 80.2 87.8 93.1 34.1 57.2 71.0 81.4
 CSVR(Ours) R3D 200 64.2 79.5 88.2 92.9 34.5 59.0 71.8 82.2
-Action Recognition.
-We firstly compare CSVR with the previous self-supervised video representation learning methods in terms of the downstream action recognition task.
-As shown in Table 7, when pretraining on UCF101, our CSVR consistently outperforms all existing methods with the same backbone.
-Moreover, when pretraining on large-scale dataset Kinetics400, our method also achieves the best results of $92.5\%$ and $64.5\%$ on UCF101 and HMDB51, respectively.
-We can also observe that the trained model by our CSVR surpasses the methods which use other modality of data like optical flow [155], multi-modal information [160].
-Besides, our CSVR needs fewer pretraining epochs than most existing methods.
-All these results demostrate that CSVR is both effective and efficient for self-supervised video representation learning on action recognition.
 Video Retrieval.
 Besides action recognition, we also conduct evaluations for the downstream video retrieval task on two human action datasets UCF101 and HMDB51.
-Table 8 shows the results of video retrieval of each method in terms of R@K, where R@K denotes the recall rate of top-K.
-All methods are pretrained on Kinetics400 with the resolution of 112$\times$112 for fair comparisons.
+Table VIII shows the results of video retrieval of each method in terms of R@K, where R@K denotes the recall rate of top-K.
+All methods are pretrained on Kinetics-400 with the resolution of 112$\times$112 for fair comparisons.
 The similar conclusion can be obtained that our CSVR achieves the best results on both UCF101 and HMDB51, showcasing its superiority for self-supervised features learning for human action understanding.
-5 Conclusion
+TABLE IX: Computing cost comparing with popular self-supervised video representation learning methods. All methods are pretrained on K400.
+{tabu}
+l—ccc—c Method Backbone GFLOPsParameters UCF Top-1
+MoCo[124] C3D 38.7 33.4M 60.5
+CoCLR[155] S3D 36.0 9.1M 87.9
+SpeedNet[98] S3D 36.5 9.1M 81.1
+$\rho$MoCo[65] SlowFast-R50 83.5 31.8M 92.8
+CSVR C3D 40.4 33.4M 92.8
+CSVR S3D 37.7 9.1M 90.2
+CSVR R(2+1)D 40.4 15.4M 92.8
+Computational cost.
+To compare computational costs, we enumerate the model parameters of some popular self-supervised video representation learning methods pretrained on Kinetics-400 in Table.
+IV-D.
+Following the common practices in [65, 168, 87], we report the parameters of each model during the inference phase.
+As shown, CSVR significantly improves accuracy in terms of the UCF101 action recognition downstream task compared to methods using the same backbone.
+Compared to MoCo[124], which also uses C3D as the backbone, our method boosts performance by over 30% on the UCF101 dataset.
+Additionally, CSVR using R(2+1)D as the backbone achieves comparable performance with $\rho$MoCo[65], which uses the heavier SlowFast [33] as the backbone.
+Noting that, CSVR has higher GFLOPs compared to the methods using the same backbone since we have an additional generative pose prediction branch.
+V Limitation
+CSVR is a self-supervised video representation learning framework with multiple branches.
+These branches independently learn dynamic and static features, which are then optimized collaboratively.
+To manage the complexity of optimizing multiple neural networks during pretraining, we opt for lightweight CNNs instead of stronger, but larger backbones like Transformers.
+Additionally, we see potential for integrating more advanced generative methods (e.g., diffusion models[171]) and contrastive learning approaches (e.g., BYOL[65]) into our framework, given that the primary goal of the dynamic and static branches is to learn comprehensive video representations.
+Exploring these enhancements further will be a focus of our future work.
+VI Conclusion
 This paper designs a self-supervised video representation learning framework named CSVR for action recognition by collaboratively training three branches: generative pose prediction branch, discriminative context matching branch and video generating branch.